OpenStudy (anonymous):

I am having a really hard time with this find dy/dx ^(6y)=cos (4x+y)

OpenStudy (anonymous):

thats supposed to be e^(6y)=cos (4x+y)

OpenStudy (anonymous):

\[you \ mean \ y \prime^{6y}= \cos(4x+y)\]

OpenStudy (anonymous):

!!!!?????

OpenStudy (anonymous):

no its e

OpenStudy (anonymous):